Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-02-12

Observing: Giacomo Terreran, Stefan Taubenberger

Support Team: S. Taubenberger, M. Kromer, T. Faran

TAT: Maria Teresa Botticella, Janet Chen


------------------------------------------------


Weather conditions


Start of night: thick clouds, humidity ~60%, wind 0 m/s

Middle of night: clear, humidity ~50%, wind 0 m/s 

End of night: clear, humidity ~30%, wind 0 m/s 



------------------------------------------------


Calibrations


Evening: -

Morning: PESSTO_SkyFF_mor_irz



------------------------------------------------


Science


-03:20 UT: open dome


-03:20 UT: data analysis

-04:15 UT: Follow-up:  OGLE15xl, griz, seeing ~0.5"

-05:00 UT: Standard: L745a, slit1'', gr11+gr16+gr13, seeing ~0.4"

-05:25 UT: Follow-up:  SN2016O, slit1'', gr13 2x2700s, seeing ~0.5" **

-07:00 UT: Follow-up:  PSN J12080010+1944267, slit1'', gr13 2x1800s, seeing ~0.7"

-08:15 UT: Follow-up:  SN2016adj, slit1'', gr11 1x900s + gr16 1x100s, seeing ~0.7"

-08:45 UT: Follow-up:  SN2016X, slit1'', gr11+gr16 600s each, seeing ~0.7"

-09:20 UT: Standard: EG274, slit1'', gr11+gr16+gr13, seeing ~0.7"


-10:05 UT: End of the night

 

**the first spectrum looks like the object moved along the slit, but overall the spectrum should be ok. However we see up to ~70% of flux loss in the red part of the second spectrum. We don't really know why.




Comments